How do I solve this equation for the value of x?
3/(x – 3) + 5/(x + 4) = (5x + 6)/(x2 + x - 12)

Answer by mananth(16946)

3/(x – 3) + 5/(x + 4) = (5x + 6)/(x2 + x - 12)
.
add the two fractions . LCM = (x-3)(x+4)
3(x+4)+5(x-3) / (x-3)(x+4)= (5x+6) / (x+4)(x-3)
.
3x+12+5x-15 = 5x+6
8x-5x=6-12+15
3x=9
x=3